d) The Profit function of a firm is of the form π = aQ2 + bQ + c if it is known that π =


31, 34, and 27 when Q = 1, 2, and 3 respectively, write down a set of three unknowns,


a, b, and c. Solve this system to find a, b, and c. hence find the profit when Q = 5. [7]


e) A real estate firm owns the Kasumba Garden Apartments which consists of 90


apartments. At K350 per month every apartment can be rented. However, for each K10


per month increase there will be two vacancies with no possibility of filling them. The


firm wants to receive K31980 per month from rents. What rent should be charged for


each apartment?

1
Expert's answer
2021-12-01T12:17:06-0500

d)


"a(1)^2 + b(1) + c=31""a(2)^2 + b(2) + c=34""a(3)^2 + b(3) + c=27"

"a + b + c=31""3a+b=3""5a+b=-7"

"a + b + c=31""3a+b=3""2a=-10"

"a=-5""b=18""c=18"


"\pi(Q)=-5Q^2+18Q+18"

"\pi (5)=-5(5)^2+18(5)+18=-17"

e) Let "R=" revenue.

If the price is "(350+10n)" then there will be"(90-2n)" apartments rented.


"(350+10n)(90-2n)=31980, 0\leq n\leq 45"

"31500+200n-20n^2=31980"

"20n^2-200n+480=0"

"n^2-10n+24=0"

"(n-4)(n-6)=0"

"n_1=4, n_2=6"

"350+10(4)=390"

"350+10(6)=410"



K390 per month or K410 per month.